# Analytics, have some humility: a statistical view of fourth-down decision making

### 11 Dec 2023 v2

@@ B.1 Modeling punt outcome
${\mathsf{PYOE}}_{n}:=\text{actual yardline after the }n^{th}\text{ punt}-{{\mathbb{E}}_{{\mathsf{Punt}}}^{(0)}}(\text{yardline prior to the }n^{th}\text{punt}).$ (B.1)
@@
${\mathsf{pq}}_{n}:=\alpha\cdot{\mathsf{pq}}_{n-1}+{\mathsf{PYOE}}_{n-1},$ (B.2)
@@
$\begin{split}{\mathbb{E}}_{\text{punt}}[\text{next yardline}]&=\ \vec{\alpha}\cdot\text{spline}\big{(}\text{yardline},\ \text{df}=4\big{)}+\ \beta_{1}\cdot{\mathsf{pq}}+\beta_{2}\cdot{\mathsf{pq}}\cdot\text{yardline}.\end{split}$ (B.3)
@@
B.2 Modeling field goal success probability
@@
${\mathsf{FGPA}}_{n}:=\mathbb{I}\left(n^{th}\text{ field goal is made}\right)-{P_{\mathsf{FG}}^{(0)}}(\text{yardline of the }n^{th}\text{ field goal}).$ (B.4)
Now we define kicker quality using Ravens’ kicker Justin Tucker for concreteness. Index all of Tucker’s field goals from 2006 to 2021 by $n$. We define Tucker’s kicker quality prior to field goal $n$ by a weighted sum of the field goal probability added in his previous kicks,
${\mathsf{kq}}_{n}:=\alpha\cdot{\mathsf{kq}}_{n-1}+{\mathsf{FGPA}}_{n-1},$ (B.5)
@@
$\begin{split}\log\bigg{(}\frac{{\mathbb{P}}(\text{make }{\mathsf{FG}})}{1-{\mathbb{P}}(\text{make }{\mathsf{FG}})}\bigg{)}=\ \vec{\alpha}\cdot\text{spline}\big{(}\log(\text{yards to go}),\ \text{df}=5\big{)}+\beta\cdot{\mathsf{kq}}.\end{split}$ (B.6)
Fitting this model on our dataset of $15,472$ observed field goals from 2006 to 2021 yields nontrivial probability predictions for extremely long field goals that have never before been made (e.g., nontrivial probability for a 73 yard field goal from the 55 yardline). To shrink these field goal probability predictions to zero, we impute 500 synthetic missed field goals, randomly distributed from the 51 to the 99 yardline, into our dataset. In Figure 12b we plot the probability of making a field goal according to our model as a function of yardline for various kicker quality values.
@@
B.3 Modeling conversion probability
@@
${\mathsf{EPA}}_{n}={\mathsf{EP}}_{n}-{\mathsf{EP}}_{n-1}.$ (B.7)
@@
$q_{n}:=\gamma_{n}\cdot\alpha\cdot q_{n-1}+{\mathsf{EPA^{(0)}}}_{n-1},$ (B.8)
where $q_{0}:=0$ and ${\mathsf{EPA^{(0)}}}_{0}:=0$. We use an exponential decay weight $\alpha$ to upweight more recent plays (Medvedovsky and Patton,, 2022). We set $\alpha=0.995$ for each of our team and player quality metrics. For instance, a play which occured 138 plays ago is weighted half as much as the previous play since $0.50.995^{138}=$ 0.5. Also, a play which occured 459 plays ago is weighted one-tenth as much as the previous play since $0.10.995^{459}=$ 0.1. Additionally, at the start of each season, we shrink by a multiplicative factor $\gamma$. So, the shrinkage weight $\gamma_{n}$ of play $n$ is $\gamma$ if this is Mahomes’ first play of the season, otherwise it is 1. We use different shrinkage values $\gamma$ for each team and player quality metric, shown in Table 3. We tuned these $\gamma$ values to optimize predictive performance on a held-out validation set. A smaller $\gamma$ reflects a quantity that is noiser from year to year. Finally, we standardize each offensive and defensive quality metric to have mean zero and standard deviation $1/2$. In Figure 13 we plot the career mean quarterback quality of each quarterback with over 1750 attempts and whose rookie season came after 2006. As expected, Patrick Mahomes has by far the highest quarterback quality. We construct the other team quality metrics described in Table 3 in a similar fashion, via Formula (B.8).
@@
Table 3: Summary of team quality metrics computed via Formula (B.8) and their start-of-season shrinkage parameters $\gamma$. Each of these metrics apply to both the offensive and defensive teams in a play, yielding eight total metrics.
Due to a small sample size of fourth-down conversion attempts, some existing models use third down as a proxy for fourth down, as they are also high pressure situations in which the offensive team attempts to gain a first down on that play (Romer,, 2006). There may be, however, a fundamental difference in conversion probability between third and fourth-down plays, perhaps due to psychological reasons or a different distribution of play calls on fourth down. Therefore, in our model selection process, we test models on a random 50% of fourth-down plays, and we train some models on a dataset consisting entirely of fourth-down plays and other models on a dataset consisting of third and fourth-down plays. We find that the parameters of our best conversion probability model borrow strength from third down plays.
@@
$\begin{split}\log\bigg{(}\frac{{\mathbb{P}}({\text{convert}})}{1-{\mathbb{P}}({\text{convert}})}\bigg{)}=&\ \vec{\alpha_{1}}\cdot\mathbb{I}\left(\text{fourth down}\right)\cdot\text{spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \vec{\alpha_{2}}\cdot\mathbb{I}\left(\text{third down}\right)\cdot\text{spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \beta_{0}+\beta_{1}\cdot{\mathsf{qbqot}}+\beta_{2}\cdot{\mathsf{oqrot}}+\beta_{3}\cdot{\mathsf{dqdtap}}+\beta_{4}\cdot{\mathsf{dqdtar}}.\end{split}$ (B.9)
@@
A more elaborate conversion probability model may adjust for yardline. In particular, it is plausible that it is more difficult to convert near each endzone, where space is more constricted. Additionally, a more fine-grained model would be continuous in yards to go rather than treating it as an integer. Of course, our model can only be as good as available data, which treats yards to go as an integer. But, anecdotally, $4^{th}$ down and inches has a significantly higher conversion probability than fourth down and one yard to go. Finally, conversion probability depends on the offensive and defensive play call. On this view, in practice it may be better for teams to input their own conversion probability estimates as they are more informed on their play calling tendecies.
Appendix C First-down win probability model details
@@
$\texttt{scoreTimeRatio}=\frac{\texttt{score\_differential}}{0.01+\texttt{game\_seconds\_remaining}}.$ (C.1)
@@
Now we compare the out-of-sample predictive performance of various ${\mathsf{WP}}$ models. Our full dataset consists of all football plays from 2006 to 2021. The dataset is clustered into games, as plays from each game share the same winning team. To keep the clustered nature of our dataset intact and to avoid data bleed, we split our dataset in half by randomly sampling 50% of all games. The first-down plays from the first 50% of these games form the hold-out test set. We test on first-down plays because fourth-down decision making relies on first-down win probability. The plays from the other 50% of these games form the training set. To tune ${\mathsf{XGBoost}}$ models, we split the training set in half by randomly sampling 50% of the games from the training set. The plays from the first 50% of these games form the ${\mathsf{XGBoost}}$ training set, and the remaining plays form the validation set for hyperparameter tuning. We then tune our ${\mathsf{XGBoost}}$ models in a similar fashion as Baldwin, (2021). We visualize the results of our model comparison in Table 4.
@@
Appendix D Simulation study details
@@
$X_{g,n+1}:=\begin{cases}X_{gn}+\xi_{gn}&\text{ if }0<X_{gn}+\xi_{gn}<L\text{ (not a TD)}\\ L/2&\text{ else,}\end{cases}$ (D.2)
@@
$S_{g,n+1}:=\begin{cases}S_{gn}+1&\text{ if }X_{gn}+\xi_{gn}=0\text{ (TD)}\\ S_{gn}-1&\text{ if }X_{gn}+\xi_{gn}=L\text{ (opp. TD)}\\ S_{gn}&\text{ else.}\end{cases}$ (D.3)
@@
$y_{gn}\equiv y_{g,N+1}:=\begin{cases}1&\text{ if }S_{g,N+1}>0\\ 0&\text{ if }S_{g,N+1}<0\\ \text{Bernoulli}(1/2)&\text{ else (overtime). }\end{cases}$ (D.4)
@@
${\mathsf{WP}}(N+1,x,s)=\begin{cases}1&\text{if }s>0\\ 1/2&\text{if }s=0\\ 0&\text{if }s<0,\\ \end{cases}$ (D.7)
@@
${\mathsf{WP}}(n-1,x,s)=\begin{cases}\frac{1}{2}{\mathsf{WP}}(n,\frac{L}{2},s+1)+\frac{1}{2}{\mathsf{WP}}(n,x+1,s)&\text{if }x=1\\ \frac{1}{2}{\mathsf{WP}}(n,x-1,s)+\frac{1}{2}{\mathsf{WP}}(n,\frac{L}{2},s-1)&\text{if }x=L-1\\ \frac{1}{2}{\mathsf{WP}}(n,x-1,s)+\frac{1}{2}{\mathsf{WP}}(n,x+1,s)&\text{else.}\end{cases}$ (D.8)
Appendix E Tuning the number of bootstrapped datasets
As discussed in Section 3.1, we use bootstrap percentage to incorporate uncertainty quantification into fourth-down decision making because it quantifies uncertainty in the decision itself. Although the strength of a decision is proportional to the ${\mathsf{boot\%}}$, we cannot necessarily rely on a non-technical football coach to fully process how ${\mathsf{boot\%}}$ maps to the strength of a decision recommendation. To faciliate easier communication to a coach, we bin decision recommendations into three buckets: confident (${\mathsf{boot\%}}\in[90\%,100\%]$), lean (${\mathsf{boot\%}}\in[75\%,90\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,75\%]$).
Since we ultimately bucket decision recommendations into three bins (see Section 4.2), we want the number of bootstrapped models $B$ to be large enough to stably categorize decisions into one of these three bins. We also want $B$ to be as small as possible in order to quickly evaluate a fourth-down recommendation during a football game. Hence we conduct a stability analysis to choose $B$.
@@ Appendix E Tuning the number of bootstrapped datasets
$\begin{cases}p_{i}^{\text{confident, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[90\%,100\%]\right),\\ p_{i}^{\text{lean, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[75\%,90\%)\right),\\ p_{i}^{\text{uncertain, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[0\%,75\%)\right),\\ p_{i}^{(B)}=\max\{p_{i}^{\text{confident, }(B)},\ p_{i}^{\text{lean, }(B)},\ p_{i}^{\text{uncertain, }(B)}\}.\end{cases}$ (E.1)
For each category, $p_{i}^{\text{category, }(B)}$ is the proportion of the $M$ bootstrap draws that play $i$ is put in that category. For our procedure with $B$ bootstrapped models to be stable, we want $p_{i}^{\text{category, }(B)}$ to be close to 1 for each category, which means that a play’s categorization is not dependent on the randomness inherent in generating $B$ bootstrapped models. Hence we want $p_{i}^{(B)}$, the maximum $p_{i}^{\text{category, }(B)}$ across the three categories, to be close to 1. Hence we want the mean across all plays $\overline{p}^{(B)}=\frac{1}{n}\sum_{i=1}^{n}p_{i}^{(B)}$ to be close to 1. In Figure 16 we see that for $B=100$, $\overline{p}^{(B)}=0.9$ and the vast majority of plays $i$ have $p_{i}^{(B)}=1$. Some plays have $p_{i}^{(B)}$ lower than 1 since they lie near the border of two categories. We believe $\overline{p}^{(B)}=0.9$ is sufficiently large and hence use $B=100$ in our fourth-down decision procedure.
@@
Appendix F Baseline coach’s decision model
To compare our decision making procedure to the decisions that actual football coaches tend to make, we model the probability that a coach chooses a decision in $\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ as a function of game-state. We use ${\mathsf{XGBoost}}$ to fit these coach probabilities. ${\mathsf{XGBoost}}$ works well here because we have $94,786$ fourth-down plays in our full dataset of plays since 1999, and each play is an independent observation of a coach’s decision. In particular, we fit these coach probabilities as a function of yardline, yards to go, game seconds remaining, score differential, point spread, and era (1999-2001, 2002-2005, 2006-2013, 2014-2017, and 2018-present). In Figure 17 we visualize these coach decision models, and the results make intuitive sense. For the most part, coaches punt deep in their own territory and kick field goals near the opponent’s endzone, except for with one and sometimes two yards to go. Also, at the end of the game, coaches’ decision making changes depending on the number of points they need to score to win the game.

### 19 Jan 2024 v4

@@ C.1 Modeling punt outcome
${\mathsf{PYOE}}_{n}:=\text{actual yardline after the }n^{th}\text{ punt}-{{\mathbb{E}}_{{\mathsf{Punt}}}^{(0)}}(\text{yardline prior to the }n^{th}\text{punt}).$ (21)
@@
${\mathsf{pq}}_{n}:=\alpha\cdot{\mathsf{pq}}_{n-1}+{\mathsf{PYOE}}_{n-1},$ (22)
@@
$\begin{split}{\mathbb{E}}_{\text{punt}}[\text{next yardline}]&=\ \vec{\alpha}\cdot\text{spline}\big{(}\text{yardline},\ \text{df}=4\big{)}+\ \beta_{1}\cdot{\mathsf{pq}}+\beta_{2}\cdot{\mathsf{pq}}\cdot\text{yardline}.\end{split}$ (23)
@@
C.2 Modeling field goal success probability
@@
${\mathsf{FGPA}}_{n}:=\mathbb{I}\left(n^{th}\text{ field goal is made}\right)-{P_{\mathsf{FG}}^{(0)}}(\text{yardline of the }n^{th}\text{ field goal}).$ (24)
Now we define kicker quality using the Ravens’ kicker Justin Tucker for concreteness. Index all of Tucker’s field goals from 2006 to 2021 by $n$. We define Tucker’s kicker quality prior to field goal $n$ by a weighted sum of the field goal probability added in his previous kicks,
${\mathsf{kq}}_{n}:=\alpha\cdot{\mathsf{kq}}_{n-1}+{\mathsf{FGPA}}_{n-1},$ (25)
@@
$\begin{split}\log\bigg{(}\frac{{\mathbb{P}}(\text{make }{\mathsf{FG}})}{1-{\mathbb{P}}(\text{make }{\mathsf{FG}})}\bigg{)}=\ \vec{\alpha}\cdot\text{spline}\big{(}\log(\text{yards to go}),\ \text{df}=5\big{)}+\beta\cdot{\mathsf{kq}}.\end{split}$ (26)
Fitting this model on our dataset of $15,472$ observed field goals from 2006 to 2021 yields nontrivial probability predictions for extremely long field goals that have never before been made (e.g., nontrivial probability for a 73-yard field goal from the 55 yardline). To shrink these field goal probability predictions to zero, we impute 500 synthetic missed field goals, randomly distributed from the 51 to the 99 yardline, into our dataset. In Figure 8b we plot the probability of making a field goal according to our model as a function of yardline for various kicker quality values.
@@
C.3 Modeling conversion probability
@@
${\mathsf{EPA}}_{n}={\mathsf{EP}}_{n}-{\mathsf{EP}}_{n-1}.$ (27)
@@
$q_{n}:=\gamma_{n}\cdot\alpha\cdot q_{n-1}+{\mathsf{EPA^{(0)}}}_{n-1},$ (28)
where $q_{0}:=0$ and ${\mathsf{EPA^{(0)}}}_{0}:=0$. We use an exponential decay weight $\alpha$ to upweight more recent plays. We set $\alpha=0.995$ for each of our team and player quality metrics. For instance, a play which occurred 138 plays ago is weighted half as much as the previous play since $0.50.995^{138}=$ 0.5. Also, a play which occurred 459 plays ago is weighted one-tenth as much as the previous play since $0.10.995^{459}=$ 0.1. Additionally, at the start of each season, we shrink by a multiplicative factor $\gamma$. So, the shrinkage weight $\gamma_{n}$ of play $n$ is $\gamma$ if this is Mahomes’ first play of the season, otherwise it is 1. We use different shrinkage values $\gamma$ for each team and player quality metric, shown in Table 4. We tuned these $\gamma$ values to optimize predictive performance on a held-out validation set. A smaller $\gamma$ reflects a quantity that is noisier from year to year. Finally, we standardize each offensive and defensive quality metric to have mean zero and standard deviation $1/2$. In Figure 9 we plot the career mean quarterback quality of each quarterback with over 1750 attempts and whose rookie season came after 2006. As expected, Patrick Mahomes has by far the highest quarterback quality. We construct the other team quality metrics described in Table 4 in a similar fashion, via Formula (28).
@@
Table 4: Summary of team quality metrics computed via Formula (28) and their start-of-season shrinkage parameters $\gamma$. Each of these metrics apply to both the offensive and defensive teams in a play, yielding eight total metrics.
Due to a small sample size of fourth-down conversion attempts, some existing models use third down as a proxy for fourth down, as they are also high-pressure situations in which the offensive team attempts to gain a first down on that play (Romer,, 2006). There may be, however, a fundamental difference in conversion probability between third and fourth-down plays, perhaps due to psychological reasons or a different distribution of play calls on fourth down. Therefore, in our model selection process, we test models on a random 50% of fourth-down plays, and we train some models on a dataset consisting entirely of fourth-down plays and other models on a dataset consisting of third and fourth-down plays. We find that the parameters of our best conversion probability model borrow strength from third-down plays.
@@
$\begin{split}\log\bigg{(}\frac{{\mathbb{P}}({\text{convert}})}{1-{\mathbb{P}}({\text{convert}})}\bigg{)}=&\ \vec{\alpha_{1}}\cdot\mathbb{I}\left(\text{fourth down}\right)\cdot\text{spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \vec{\alpha_{2}}\cdot\mathbb{I}\left(\text{third down}\right)\cdot\text{spline}\big{(}\log(\text{yards to go}),\ \text{df}=4\big{)}\\ +&\ \beta_{0}+\beta_{1}\cdot{\mathsf{qbqot}}+\beta_{2}\cdot{\mathsf{oqrot}}+\beta_{3}\cdot{\mathsf{dqdtap}}+\beta_{4}\cdot{\mathsf{dqdtar}}.\end{split}$ (29)
@@
A more elaborate conversion probability model may adjust for yardline. In particular, it is plausible that it is more difficult to convert near each endzone, where space is more constricted. Additionally, a more fine-grained model would be continuous in yards to go rather than treating it as an integer. Of course, our model can only be as good as available data, which treats yards to go as an integer. But, anecdotally, fourth down and inches has a significantly higher conversion probability than fourth down and one yard to go. Finally, conversion probability depends on the offensive and defensive play call. On this view, in practice it may be better for teams to input their own conversion probability estimates as they are more informed on their play calling tendencies.
Appendix D First-down win probability model details
@@
$\texttt{scoreTimeRatio}=\frac{\texttt{score\_differential}}{0.01+\texttt{game\_seconds\_remaining}}.$ (30)
@@
Now we compare the out-of-sample predictive performance of various ${\mathsf{WP}}$ models. Our full dataset consists of all football plays from 2006 to 2021. The dataset is clustered into games, as plays from each game share the same winning team. To keep the clustered nature of our dataset intact and to avoid data bleed, we split our dataset in half by randomly sampling 50% of all games. The first-down plays from the first 50% of these games form the hold-out test set. We test on first-down plays because fourth-down decision making relies on first-down win probability. The plays from the other 50% of these games form the training set. To tune ${\mathsf{XGBoost}}$ models, we split the training set in half by randomly sampling 50% of the games from the training set. The plays from the first 50% of these games form the ${\mathsf{XGBoost}}$ training set, and the remaining plays form the validation set for hyper-parameter tuning. We then tune our ${\mathsf{XGBoost}}$ models in a similar fashion as Baldwin, (2021). We visualize the results of our model comparison in Table 5.
@@
Appendix E Simulation study details
@@
$X_{g,n+1}:=\begin{cases}X_{gn}+\xi_{gn}&\text{ if }0<X_{gn}+\xi_{gn}<L\text{ (not a TD)}\\ L/2&\text{ else,}\end{cases}$ (32)
@@
$S_{g,n+1}:=\begin{cases}S_{gn}+1&\text{ if }X_{gn}+\xi_{gn}=0\text{ (TD)}\\ S_{gn}-1&\text{ if }X_{gn}+\xi_{gn}=L\text{ (opp. TD)}\\ S_{gn}&\text{ else.}\end{cases}$ (33)
@@
$y_{gn}\equiv y_{g,N+1}:=\begin{cases}1&\text{ if }S_{g,N+1}>0\\ 0&\text{ if }S_{g,N+1}<0\\ \text{Bernoulli}(1/2)&\text{ else (overtime). }\end{cases}$ (34)
@@
${\mathsf{WP}}(N+1,x,s)=\begin{cases}1&\text{if }s>0\\ 1/2&\text{if }s=0\\ 0&\text{if }s<0,\\ \end{cases}$ (37)
@@
${\mathsf{WP}}(n-1,x,s)=\begin{cases}\frac{1}{2}{\mathsf{WP}}(n,\frac{L}{2},s+1)+\frac{1}{2}{\mathsf{WP}}(n,x+1,s)&\text{if }x=1\\ \frac{1}{2}{\mathsf{WP}}(n,x-1,s)+\frac{1}{2}{\mathsf{WP}}(n,\frac{L}{2},s-1)&\text{if }x=L-1\\ \frac{1}{2}{\mathsf{WP}}(n,x-1,s)+\frac{1}{2}{\mathsf{WP}}(n,x+1,s)&\text{else.}\end{cases}$ (38)
Appendix F Fitting our fourth-down decision procedure
In this Section we describe in detail our fourth-down decision procedure, which incorporates uncertainty quantification.
We begin by fitting the traditional fourth-down decision procedure. We fit our first-down win probability model, an ${\mathsf{XGBoost}}$ model fit from first-down plays, to our dataset (recall Section 2.4 and Appendix D). We also fit our decision transition models from our dataset (recall the punt outcome model, field goal success probability model, and conversion probability model, described in Appendix C). We then knit these models together to estimate fourth-down win probability for each decision ${\mathsf{Go}}$, ${\mathsf{FG}}$, and ${\mathsf{Punt}}$ (described in Section 2 and Appendix B). The estimated optimal decision is the one that maximizes win probability.
We use bootstrapping to quantify uncertainty in our fourth-down decision recommendations. We generate $B=100$ bootstrapped datasets using a randomized cluster bootstrap as follows (we justify our choice of $B=100$ in Appendix G). Let $G$ denote the number of games in our observed play-by-play dataset. To generate a bootstrapped dataset we re-sample $G\cdot\phi$ games uniformly with replacement. We use $\phi=0.5$ as justified by the simulation study from Section 3.1 in order to achieve adequate marginal coverage. Within each game $g$, which contains $N_{g}$ drives, we re-sample $N_{g}$ drives uniformly with replacement. We re-sample drives with replacement, rather than plays with replacement as we did in the simulation study, since it provides a more realistic re-sampling of real football data.
From each of the $B$ bootstrapped datasets we fit a first-down win probability model. We knit each bootstrapped first-down win probability model together with the decision transition models fit from observed data to form $B$ bootstrapped fourth-down decision models. For each game-state, from the $B$ bagged decisions we calculate bootstrap percentage and form a $90\%$ confidence interval on the estimated gain in win probability.
Appendix G Tuning the number of bootstrapped datasets
As discussed in Section 3.2, we use bootstrap percentage to incorporate uncertainty quantification into fourth-down decision making because it quantifies uncertainty in the decision itself. Although the strength of a decision is proportional to the ${\mathsf{boot\%}}$, we cannot necessarily rely on a non-technical football coach to fully process how ${\mathsf{boot\%}}$ maps to the strength of a decision recommendation. To facilitate easier communication to a coach, we bin decision recommendations into three buckets: confident (${\mathsf{boot\%}}\in[90\%,100\%]$), lean (${\mathsf{boot\%}}\in[75\%,90\%)$), and uncertain (${\mathsf{boot\%}}\in[0\%,75\%]$).
Since we ultimately bucket decision recommendations into three bins (see Section 4.2), we want the number of bootstrapped models $B$ to be large enough to stably categorize decisions into one of these three bins. We also want $B$ to be as small as possible in order to quickly evaluate a fourth-down recommendation during a football game. Hence, we conduct a stability analysis to choose $B$.
@@ Appendix G Tuning the number of bootstrapped datasets
$\begin{cases}p_{i}^{\text{confident, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[90\%,100\%]\right),\\ p_{i}^{\text{lean, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[75\%,90\%)\right),\\ p_{i}^{\text{uncertain, }(B)}=\frac{1}{M}\sum_{m=1}^{M}\mathbb{I}\left({\mathsf{boot\%}}_{im}^{(B)}\in[0\%,75\%)\right),\\ p_{i}^{(B)}=\max\{p_{i}^{\text{confident, }(B)},\ p_{i}^{\text{lean, }(B)},\ p_{i}^{\text{uncertain, }(B)}\}.\end{cases}$ (39)
For each category, $p_{i}^{\text{category, }(B)}$ is the proportion of the $M$ bootstrap draws that play $i$ is put in that category. For our procedure with $B$ bootstrapped models to be stable, we want $p_{i}^{\text{category, }(B)}$ to be close to 1 for each category, which means that a play’s categorization is not dependent on the randomness inherent in generating $B$ bootstrapped models. Hence, we want $p_{i}^{(B)}$, the maximum $p_{i}^{\text{category, }(B)}$ across the three categories, to be close to 1. Hence, we want the mean across all plays $\overline{p}^{(B)}=\frac{1}{n}\sum_{i=1}^{n}p_{i}^{(B)}$ to be close to 1. In Figure 12 we see that for $B=100$, $\overline{p}^{(B)}=0.9$ and the vast majority of plays $i$ have $p_{i}^{(B)}=1$. Some plays have $p_{i}^{(B)}$ lower than 1 since they lie near the border of two categories. We believe $\overline{p}^{(B)}=0.9$ is sufficiently large and hence use $B=100$ in our fourth-down decision procedure.
@@
Appendix H Baseline coaches’ decision model
To compare our decision-making procedure to the decisions that actual football coaches tend to make, we model the probability that a coach chooses a decision in $\{{\mathsf{Go}},{\mathsf{FG}},{\mathsf{Punt}}\}$ as a function of game-state. We use ${\mathsf{XGBoost}}$ to fit these coach probabilities. ${\mathsf{XGBoost}}$ works well here because we have $94,786$ fourth-down plays in our full dataset of plays since 1999, and each play is an independent observation of a coach’s decision. In particular, we fit these coach probabilities as a function of yardline, yards to go, game seconds remaining, score differential, point spread, and era (1999-2001, 2002-2005, 2006-2013, 2014-2017, and 2018-present). In Figure 13 we visualize these coach decision models, and the results make intuitive sense. For the most part, coaches punt deep in their own territory and kick field goals near the opponent’s endzone, except for with one and sometimes two yards to go. Also, at the end of the game, coaches’ decision-making changes depending on the number of points they need to score to win the game.